OpenStudy (anonymous):

Sketch the following regions and state the interior and the closure: a) |z-2+i|≤1 b) Im(z)>1

OpenStudy (anonymous):

a) z=x+iy so |x+iy-2+i|-> |(x-2)+i(y+1)|≤1 So (x-2)2+(y+1)2≤1 So it would just be a circle on the real plane? And the interior would be the equation with < instead of ≤ right? I'm not sure how to write the closure though. b)Im(z)= y so it would be a straight line at y=1 on the real plane and there is no interior or closure since it is an open set right?

OpenStudy (anonymous):

@satellite73 @phi @UnkleRhaukus

OpenStudy (anonymous):

The closure of b) is the set \[ \{ z \in C | Im(z) \ge 1\} \] So it contains the line y=1 and anything above it in the xy-plane

OpenStudy (anonymous):

@eliassaab isn't that still open though? If it's everything above y=1 then there is no set closure??

OpenStudy (unklerhaukus):

right, the solution set for b) is not closed for example z_1 = 2i , z_2 = 3i . z-1-z_3 = -i which is not in the solution set

OpenStudy (anonymous):

I'm a bit confused as to what it does have. Does that mean the interior is Im(z)>1? @UnkleRhaukus

OpenStudy (unklerhaukus):

i think so, but i dont really understand the concept of interior